AubreeTheVaimpier:

Can someone plz help me add words to this? (I'll also have the question with it) You need to download 3 games for your math class. Each game is the same price. Since this is your first time downloading a game from the site, you are also charged a one-time $0.50 access fee. You are charged a total of $4.35. What is the cost of each game? Show your work and round your answer to the nearest cent. 4.35-.50= 3.85 divide that by 3= 1.28$ per game

AubreeTheVaimpier:

@jhonyy9

AubreeTheVaimpier:

@jhonyy9

jhonyy9:

AZ here pls - ty

Astrid1:

@AZ

AZ:

Hello

AZ:

We have to come up with an equation and solve it So we're buying 3 games Each game costs `x` dollars There's also a one time charge of 0.50 The total price comes out to be 4.35 Do you have any idea how to write this all as one equation?

AubreeTheVaimpier:

3(x+0.50)=4.35 or \[4.35-.50=3.85\div3=x\]?

AZ:

Soooooooooo

AZ:

Your work is correct but the first equation isn't

AZ:

That's because when you have 3(x+0.5) that means you're multiplying the 3 inside it should just be 3x + 0.5

AubreeTheVaimpier:

oh

AZ:

If you have 3(x+0.5) that means 3x + 3*0.5 or 3x + 1.5 but the 0.5 was a one time fee so we can't multiply it by 3 Did you understand? Or would you like me to explain it more

AubreeTheVaimpier:

A little more explain

AZ:

so you have 3 games, right? Each game costs x dollars so how much does 3 games cost?

AZ:

as an example: you have 3 apples each apple costs 5 dollars how much does 3 apples cost? it costs 3 * 5 so similarly 3 games each costs x dollars how much does 3 games cost?

AubreeTheVaimpier:

I have how much each game cost already. I want to have some words with it like this one, since I sent about 12 text messages a day and my older sister sends more than me and together we send a total of about 26 messages per day, you can do 26 (total messages) - 12 (my text messages) = 14 (my sister’s text messages).

AZ:

Is that like another question? I think I'm missing something here. What is the instructions for your question/assignment?

AubreeTheVaimpier:

Here are the directions

1 attachment
AZ:

So did you come up with the question yourself or are you just supposed to write a really good answer to the question they gave you?

AubreeTheVaimpier:

Heres the question

1 attachment
AZ:

Well we've already answered it all In math, you don't really need to add too many words to your explanation :)

AubreeTheVaimpier:

wait we did

AZ:

so first, what is the equation? Scroll up- remember I corrected it for you?

AubreeTheVaimpier:

This

1 attachment
AZ:

yes and obviously don't forget the equal to 4.35

AubreeTheVaimpier:

So just put 3x+0.5=4.35

AZ:

That's the equation! And then you solved it correctly First you have to subtract 0.5 on both sides And then you divide 3 on both sides and you've solved for x

AZ:

You might want to write it out line by line instead of it all as one long line